# MetaTool: Facilitating Large Language Models to Master Tools with Meta-task Augmentation

### v1

@@ Abstract
Utilizing complex tools with Large Language Models (LLMs) is a critical component for grounding AI agents in various real-world scenarios. The core challenge of manipulating tools lies in understanding their usage and functionality. The prevailing approach involves few-shot prompting with demonstrations or fine-tuning on expert trajectories. However, for complex tools and tasks, mere in-context demonstrations may fail to cover sufficient knowledge. Training-based methods are also constrained by the high cost of dataset construction and limited generalizability. In this paper, we introduce a new tool learning methodology (MetaTool) that is generalizable for mastering any reusable toolset. Our approach includes a self-supervised data augmentation technique that enables LLMs to gain a comprehensive understanding of various tools, thereby improving their ability to complete tasks effectively. We develop a series of meta-tasks that involve predicting masked factors of tool execution. These self-supervised tasks enable the automatic generation of high-quality QA data concerning tool comprehension. By incorporating meta-task data into the instruction tuning process, the proposed MetaTool model achieves significant superiority to open-source models and is comparable to GPT-4/GPT-3.5 on multiple tool-oriented tasks.
@@ 1 Introduction
Distinguished from other species, a critical characteristic of human beings’ advanced intelligence is the use of complex tools, which expands the frontiers neural intelligence can reach. With the advent of powerful foundation models (e.g. large language models, multi-modal models), AI has the potential to solve complex tasks as a general agent, equipped with the abilities to make long-term plans, use external tools, reflect on its own behavior, etc. Using tools crucially endows LLMs the power from external mechanisms and to exert effect on a larger scale.
Existing tool learning research majorly falls into two paradigms: tool-augmented learning and tool-oriented learning Qin et al. (2023b). The former aims at augmenting the model with complementary resources (e.g. retriever, search engine), and the latter focuses on achieving certain task objectives with tools (e.g. web navigationRawles et al. (2023); Hong et al. (2024), embodied manipulationChi et al. (2023)). While augmenting LLMs with tools requires appropriate tool selection, tool-oriented tasks raise more challenges in tool manipulation given that the orientation is tool output and state change. This work focuses on learning tool manipulation in the tool-oriented paradigm with large language models.
To utilize tools with LLMs, a mainstream way is to provide the ”cookbook” of tools with zero-shot prompting or demonstrations of tool usage with few-shot promptingXu et al. (2023); Brown et al. (2020). It may work on simple tool sets, however, for complex tools like software or machines, demonstrations can not exhaust all scenarios, and manuals are also limited in length. Ultimately, it’s impractical to expect a system to be intelligent enough to master any tool without experience of using it. Besides, prior training-based methods mainly adopt supervised fine-tuning with annotated solutions on the basis of pre-trained LLMs Qin et al. (2023c); Patil et al. (2023). Regardless of the difficulties of annotating the optimal actions for complex tasks, training on a limited amount of data is prone to overfitting action patterns. Without truly understanding the dynamics of tool execution, it’s hard to generalize to diverse task scenarios through flexible tool manipulation. For instance in Figure 1, understanding the usage or functionality of hammers (e.g. nailing, smashing) enables the robot to build a cabin better and generalize the skills to chair or table construction.
@@
Figure 1: Paradigm comparison between existing tool learning methods and proposed meta-task augmentation.
Towards the issues above, our insight is that generalizable tool manipulation should be achieved on the foundation of comprehensive tool understanding, which is learnable with a practical amount of data. In this paper, we propose a data augmentation method (MetaTool) that enhances LLM’s understanding of an external toolset and boosts the learning of tool-oriented tasks. Given a callable toolset (e.g. APIs, programs), a meta-set consisting of question-answering data of 6 meta-tasks is constructed by calling the tools in a self-supervised way. The meta-tasks are designed concerning the causality of the toolset as an autonomous system and its functionality as a function. Then we augment the solution data of tool-oriented tasks with the meta-set to fine-tune the pre-trained LLM. Evaluated on three tool-oriented tasks, our method significantly improves the success rate (+22.7%) of the open-source LLM (e.g. LLaMA-3) and is competitive with GPT-4/3.5-turbo. Moreover, we also explore the mechanism of enhancing tool manipulation capability with meta-task data. The overall contribution can be summarized in three-folds:
@@
  • We introduce a new tool learning paradigm that facilitates the task performance of LLMs with task-agnostic tool understanding.
@@
  • We propose an integral set of self-supervised tasks that dissect the tool execution process and enable efficient data generation and augmentation.
  • Extensive evaluation on tool-oriented tasks verifies that MetaTool significantly enhances open-source LLMs compared with conventional instruction tuning methods.
2 Related Works
2.1 Tool learning
Recent studies have shed light on the potential of utilizing tools to augment LLMs with external factual knowledge Qin et al. (2023a); Nakano et al. (2021); Song et al. (2023); Hao et al. (2024); Shen et al. (2024); Gao et al. (2023); Wu et al. (2023); Qian et al. (2023); Zhuang et al. (2024); Schick et al. (2024) and complete tasks in complex environments Gupta & Kembhavi (2023). With the burgeoning intelligence in reasoning and perception, LLMs’ tool-use capability can be widely applied in the automation of various domains including Embodied AI Wang et al. (2024c; b), web manipulation Rawles et al. (2023); Hong et al. (2024); Yang et al. (2023); Deng et al. (2024); He et al. (2024); Zhou et al. (2023), and image/video editing Wang et al. (2024a); Argaw et al. (2022); Hang et al. (2024); Fu et al. (2023). Effectively mastering complex tools challenges the model to comprehend the precondition and potential outcome of using tools. In this paper, we aim to facilitate LLMs for tool-oriented tasks by learning robust tool understanding.
2.2 Tool understanding
3 Method
@@
In this section, we first formalize the tool-oriented task with a close toolset. Then we define five general meta-tasks that are key to tool understanding and show how to generate datasets of them in an integral self-supervised way. In the end, we describe several training schemes to augment the tool-oriented training with meta-tasks data.
3.1 Problem Formalization
@@
A tool task can be defined as a tuple $\left\langle\mathcal{S},\mathcal{A},\mathcal{T},g\right\rangle$, where $\mathcal{S},\mathcal{A},\mathcal{T}$ is the state space, action space, and toolset, and $g$ is the goal state of the task. Toolset $\mathcal{T}=\{t\}_{N}$ consists of $N$ tools, each as a state transition function $s^{\prime}=t(s,\theta)$ that formalizes the outcome of state change when feeding the input parameters $\theta$ into the tool. An action $a=\left\langle t,\theta\right\rangle\in\mathcal{A}$ specifies the tool and its input. As an autonomous agent, an LLM should iteratively respond with actions and inputs according to the state until it reaches the goal. Broadly, when the tools can not alter any external state, tool output like retrieval results can be regarded as the state thus the goal state is the desired information.
@@
3.2 Self-supervised Meta-tasks for Tool Understanding
Meta-task definition. We first generate single-step tool execution data $\mathcal{D}=\{s,a,s^{\prime}\}$ by stochastically sampling initial state $s\in\mathcal{S}$ and action $a\in\mathcal{A}$, and obtaining the tool output $s^{\prime}$. Five surrogate tasks (meta-tasks) are designed based on the unsupervised dataset $\mathcal{D}$. Basically, the model is required to predict masked factors of the execution process in line with the idea of Masked Autoencoder (MAE) He et al. (2022). We define the meta-tasks as below:
@@ 3.2 Self-supervised Meta-tasks for Tool Understanding
  • Input Boundary: The model determines whether an action can be successfully executed given the current state: $P(\mathbbm{1}_{s^{\prime}\neq s}|a,s)$.
@@
Effect, decision-making, reversion meta-tasks emphasize the causality of a tool, regarding the action as the intervention to the state Pearl (2009); Pearl & Mackenzie (2018) and the outcome as the causal effect is determined by the tool mechanism. On top of that, counterfact task is the composition of reversion and effect, further imagining the outcome altered from the fact in effect task and raising higher requirements for causal reasoning Bareinboim et al. (2015); Zhang & Bareinboim (2016). When implemented as APIs, tools may receive non-executable inputs and result in ineffective outcomes. Thus the input and output domains are also unique features of a tool as a function. We consider Input boundary meta-task that emphasizes the tool affordance that refers to what actions can be executed considering the situation and the precondition. Output boundary meta-task emphasizes the functionality of tools, that is, what goals can and cannot be achieved given the current state.
Figure 3: Demonstrations of meta-tasks and tool-oriented solution data of SAW task (refer to section 4.1). We diversify the questions and annotated thoughts in the datasets with multiple templates generated by GPT-4.
@@
Metaset construction. Based on the single-step data $\mathcal{D}$, datasets of meta-tasks (referred to as metasets) are constructed by automatically generating question-answering pairs, showcased in Figure 2. For each sample and each meta-task, we insert the variables of states and actions into 5 templates (diversified with GPT-4) to obtain diverse QA data.
3.3 Meta-task Augmentation
@@
With the data of meta-tasks, we explore several manners to augment the tool manipulation ability to achieve task goals: 1) In-context learning: To enhance the tool understanding of LLMs in a training-free way, we incorporate several demonstrations of each meta-tasks in the system prompt. Existing works may include demonstrations in the tool cookbook, yet not in a systematic way. 2) Self-supervised Learning: Since we aim to build the model’s tool understanding as the foundation of tool-oriented learning, an intuitive manner is to train the LLM first on the metasets as the surrogate tasks and then on the solution data of tool-oriented tasks. In order to maintain the general ability of the model in the first stage, only the parameters of the query and value projection layers of the Transformer are updated instead of full-parameter training. We also propose to train the model on metasets separately to build the tool-use ability step by step. 3) Data augmentation: We also utilize the metasets as the augmented data of conventional instruction tuning methods that the metasets are mixed with solution data and the model is trained uniformly. The model trained on the mixed data is referred to as MetaTool.
@@ 4 Experiments
4.1 Task Setup
To evaluate LLMs’ ability on tool-oriented tasks, we develop 3 tasks emphasizing closed toolset manipulation rather than long-term planning and reflection which are also crucial abilities for AI agents. The key challenge of these tasks is effectively utilizing tools to achieve the goal, which requires the model to understand the rules (preconditions) and the mechanisms of the toolset. The task definition and dataset construction are elaborated below.
SpellAnyWord (SAW). In this task, the agent needs to sequentially construct a string that contains the target string as a continuous substring. The initial state of the task is a void string. Two non-degradable tools (functions) are avaliable: Add: to add two adjacent letters in the alphabet to the end of the current string. The tool input $\theta$ should be the preceding letter (e.g. passing ’a’ to Add on current string ” will result in ’ab’). Swap: to swap the position of two adjacent letters in the current string. The input should be the preceding letter (e.g. passing ’a’ to Swap on ’ab’ will result in ’ba’). An example task: The target string is ’any’. A successful action sequence can be [Add(’a’), Add(’n’), Add(’y’), Swap(’a’), Add(’o’)], which will result in a state sequence [’ab’, ’abno’, ’abnoyz’, ’banoyz’, ’banyoz’] and the final string ’banyoz’ has ’any’ as a substring.
@@ 4.1 Task Setup
Datasets collection. For the SAW task, we randomly sample 2k target strings (from 2 letters to 10 letters) as task goals. We modify the BW and LOG tasks from the prior LLM benchmark Valmeekam et al. (2024) into the tool-use version, thus 2k goals for each task are adopted following the original configuration. Optimal action sequences are obtained with heuristic strategy as the solution data. For each annotated action, we generate a thought with one of 5 templates (diversified with GPT-4). The thoughts analyze the situation and what to do following ReACT Yao et al. (2022) to leverage the model’s reasoning ability. Thus the solution contains a sequence of thought-tool-input tuples. Besides the 3 tasks we define, MetaTool can be easily generalized to other tool-oriented tasks by writing the system prompt and developing the tool functions or APIs as the external environment. The meta-task data can then be generated automatically.
4.2 Implementation Details
Our model is fine-tuned based on LLaMA3-8b-instruct AI@Meta (2024) with parameter-efficient fine-tuning method Qlora Dettmers et al. (2024) on 8 A100 GPUs. We utilize the instruction tuning version of LLaMA3 since comprehending tool-oriented tasks with specific objectives is the basis of tool understanding and manipulation. For tool-oriented task training, we construct instruction-solution pairs from the task goals and annotated solutions. For meta-task training, we formulate the task objectives and the tool execution outcome as question-answering pairs. For each task, we train the model on 10k meta-task data and 10k solution data for 3 epochs with AdamW optimizer and the learning rate of 2e-4. The models are tested in a simulated environment that receives the action of using a tool and returns the outcome and current state. We evaluated the model performance on 100 unseen cases of each three tasks.

### v2

@@ Abstract
Utilizing tools with Large Language Models (LLMs) is essential for grounding AI agents in real-world applications. The prevailing approach involves few-shot prompting with demonstrations or fine-tuning with expert annotations. However, mere in-context demonstrations may fail to cover sufficient knowledge for complex tools and tasks. Training on solution paths is also hindered by the high cost of expert annotations and generalizing to new tools. A core challenge of generalizable tool use lies in understanding the “meta”, or fundamental natures of tools that are transferable across tasks, such as causality and constraints. In this paper, we present MetaTool, a novel tool learning methodology designed to generalize across any reusable toolset. Our approach incorporates a self-supervised augmentation technique derived from a series of meta-tasks. This involves predicting masked elements in the tool execution process. The self-supervised procedure enables scalable generation of high-quality QA data, which is handy for supervising tool understanding. By incorporating meta-task data into task-oriented training, our method significantly enhances the performance of open-source LLMs, achieving results comparable to ChatGPT in both tool-based planning and chatting scenarios. Through large-scale instruction tuning, the MetaTool model demonstrates impressive zero-shot generalizability on new tasks.
@@ 1 Introduction
Distinguished from other species, a unique characteristic of human advanced intelligence is using complex tools, which expands the frontiers neural intelligence can reach. With the advent of powerful foundation models, AI has the potential to solve complex tasks with these external mechanisms. LLMs have been majorly oriented towards either tool-augmented chatbots equipped with retrievers and search engines, or tool-oriented agents (e.g. web navigation Rawles et al. (2023); Hong et al. (2024), embodied manipulation Chi et al. (2023)) that achieve task objectives through tool output Qin et al. (2023b). While the former emphasizes generalizing to various tools, the latter focuses on complex tools and scenarios.
To efficiently integrate LLMs with tools, a mainstream way relies on in-context learning (ICL). The model is provided with the “cookbook” of tools in zero-shot prompting or demonstrations in few-shot prompting Xu et al. (2023); Paranjape et al. (2023); Brown et al. (2020). It may work well on simple tools with frameworks like LangChain Chase (2022). However, for complex tasks with sophisticated tools, in-context learning is limited that demonstrations can not exhaust all scenarios, and manuals are also limited in length. Ultimately, it’s impractical to expect LLMs to be intelligent enough to master any tool without the experience of using it. On the other side, training-based methods Qin et al. (2023c); Patil et al. (2023); Dubey et al. (2024) mainly adopt supervised fine-tuning with annotated expert solutions. Despite the difficulties in scaling up the optimal annotation, supervision with task solutions has limitations. Task-agnostic knowledge of tools can be neglected, which hinders the generalization to diverse scenarios or new tools. Self-play training methods like Toolformer Schick et al. (2024) and TALM Parisi et al. (2022) integrate the inference process with self-supervised tool calling data. Although such manner maintains the generality of tool calling, it’s constrained in question-answering scenarios.
@@
Figure 1: Paradigm comparison between existing tool learning methods and proposed meta-task augmentation. While the prevailing methods are limited in generalizing to complex scenarios or new tools, MetaTool enables gaining transferable tool understanding from task-agnostic knowledge.
Empirically, human learners get familiar with tools such as hammers (e.g., for nailing and smashing) before engaging in complex construction. Generalizable tool use should be achieved based on the fundamental understanding of tools themselves that holds stable for different objectives, namely task-agnostic (illustrated in Figure 1). Naturally the formation of tool understanding can be disentangled from the learning of task solving. In this paper, we introduce MetaTool, a general methodology that enables both complex tool mastery and unseen tool generalization on top of task-agnostic tool understanding. We design a set of meta-tasks inquiring about the causality of the toolset as an autonomous system and its functionality as a function. Given a callable toolset (e.g. APIs, functions), meta-task data is constructed in a scalable self-supervised way based on unsupervised or self-play tool executions. Augmenting task-oriented training with meta-task data, LLMs learn to solve problems while deepening the mastery of tools. We conduct experiments on both complex tool-oriented tasks and tool-augmented benchmarks, demonstrating that MetaTool significantly exceeds models trained merely on annotated solutions in both worlds and is competitive with the latest LLMs (ChatGPT) with the size of 8B. The overall contribution can be summarized in three-folds:
@@
  • We introduce a new tool learning method that facilitates LLMs to master tools with task-agnostic tool understanding.
@@
  • We propose an integral set of self-supervised meta-tasks that dissect the tool execution process. Meta-tasks enable expert-free data generation and augmentation across tool-augmented and tool-oriented scenarios.
  • Extensive evaluation on both tool-oriented tasks and tool-augmented benchmarks demonstrates the effectiveness and generality of MetaTool, narrowing the gap between open-source models and state-of-the-art LLMs.
2 Approach
@@
In this section, we first formalize the task of using a close toolset and define 6 general meta-tasks that are key to tool understanding. Then we show how to construct datasets in an integral self-supervised way covering different scenarios. In the end, we describe several schemes to augment tool learning with meta-tasks.
2.1 Self-supervised Meta-tasks for Tool Understanding
@@
Problem formalization. A tool-use task can be generally defined as a Markovian tuple $\left\langle\mathcal{S},\mathcal{A},\mathcal{T},g\right\rangle$, where $\mathcal{S},\mathcal{A},\mathcal{T}$ is the state space, action space, and toolset, and $g$ is the goal state of the task. Toolset $\mathcal{T}=\{t\}_{N}$ consists of $N$ tools, each as a state transition function $s^{\prime}=t(s,\theta)$ that formalizes the outcome of state change when feeding the input parameters $\theta$ into the tool. An action $a=\left\langle t,\theta\right\rangle\in\mathcal{A}$ specifies the tool and its input. As an autonomous agent, an LLM should iteratively respond with tool calling and inputs according to the state until it reaches the goal. Broadly, when the tools can not alter any external state, tool output like retrieval results can be regarded as the state, and the desired information is the goal $g$.
@@
Meta-task definition. We begin with single-step tool execution $\mathcal{D}=\{s,a,s^{\prime}\}$, peeling off the task goal $g$ context, which results as unsupervised data. Six surrogate tasks (meta-tasks) are designed based on the dataset $\mathcal{D}$. Basically, the model is required to predict masked elements of the execution process. It aligns with the idea of masked language models such as Bert Devlin (2018), but in a broader and structural granularity. We define the meta-tasks as below (Figure 2):
@@
  • Input Boundary: The model determines whether an action can be successfully executed, namely whether the action falls in the valid action space, given the current state: $P(\mathbbm{1}_{s^{\prime}\neq s}|a,s)$.
@@
Effect, decision-making, reversion meta-tasks emphasize the causality of a tool, regarding the action as the intervention to the state Pearl (2009); Pearl & Mackenzie (2018) and the outcome as the causal effect is determined by the tool mechanism. On top of that, counterfact task is the composition of reversion and effect, further imagining the outcome altered from the fact in effect task. This meta-task raises higher requirements on counterfactual reasoning Bareinboim et al. (2015); Zhang & Bareinboim (2016), an advanced form of causal reasoning that humans use to contemplate ’what if’. Moreover, tools implemented as APIs may receive non-executable inputs and result in ineffective outcomes. Thus the input and output domains are also unique features of a tool as a function. We consider input boundary meta-task that emphasizes the tool affordance that refers to what actions can be executed considering the situation and the precondition. output boundary meta-task emphasizes the functionality of tools, that is, what goals can and cannot be achieved given the current state.
Figure 3: Two-step approach to construct metaset. We illustrate two examplified processes of both tool-oriented and tool-augmented scenarios, which don’t require any expert annotation. $x_{i}^{D},y_{i}^{D}$ denotes the $i$-th question-answer pair of decision-making meta-task, et cetera.
2.2 Metaset Construction
Based on the definitions, the dataset of meta-tasks (referred to as metaset) is generated as question-answering pairs to maintain the conversational skills of LLMs. To answer the meta-task questions, the trained model needs to understand the toolset mechanisms from the corresponding aspects. Given a set of reusable and callable tools $T$, the metaset =$\mathcal{M}=\{x_{n}^{m},y_{n}^{m}\}_{n=1}^{N}$ can be constructed in two steps, as illustrated in Figure 3, where $x_{n}^{m},y_{n}^{m}$ is the $n$-th question-answer pair of meta-task $m$. First, we generate tool execution data $\mathcal{D}$ with the toolset. For a limited amount of tools and state space, stochastic sampling can be applied to initialize state $s\sim P(\mathcal{S})$ and action $a\sim P(\mathcal{A})$, and obtain the tool output $s^{\prime}$. For large toolsets and diverse task scenarios that are hard to enumerate, we incorporate LLMs with self-play or tree search techniques to reduce redundant trials $a,s^{\prime}\sim P(\mathrm{LLM}(s,g))$. Note that the LLM does not need to be proficient in tool tasks, as the execution data $\mathcal{D}$ is irrelevant to the task performance. Non-executable actions also contain valuable knowledge and can be transferred as invalid samples in input boundary meta-task. The tool execution data should be sufficient to cover the various scenarios that may arise during the tasks.
@@ 2.2 Metaset Construction
Second, for the $n$-th sample and meta-task $m$, we insert the variables of states and actions into $K$ sets of templates (diversified with GPT-4) to obtain diverse QA data.
$(x_{n}^{m},y_{n}^{m})=\mathrm{Mask}(a_{n},s_{n},s^{\prime}_{n},\Phi_{k}^{m})$ (1)
, where $\Phi_{k}^{m}$ is the sampled template for meta-task $m$. Particularly, in tool-augmented scenarios, predicting the output of tools such as QA systems can be impractical. Nonetheless, the retrieval result itself is inherently meaningless (e.g., ’sunny’); however, it gains complete significance when combined with the context of tool calling (e.g., ’the weather in New York is sunny’), as showcased in Figure 3. Thus we modify the context into a more informative state in such scenarios by prompting LLMs $s^{*}_{n}=\mathrm{LLM}(s^{\prime}_{n},a_{n},t)$, which is trivial for most language models. At last, we arrange multiple metaset pairs of the same toolset into multi-turn QA data as the metaset $\mathcal{M}$, in order to maintain multi-turn dialogue capacity.
2.3 Tool Learning Augmented with Meta-tasks
@@
By incorporating meta-task data, we explore several approaches to augment the tool learning for the purpose of achieving task objectives: a) In-context learning: We randomly sample several demonstrations of each meta-task and add them to the system prompt to facilitate tool understanding in a training-free manner. Such task-agnostic knowledge includes the interpretation of rules, supplementing the solution demonstrations. b) 2-stage learning: Since we aim to build the model’s tool understanding as the foundation of tool-oriented learning, an intuitive idea is to train the LLM first on the metasets as the surrogate tasks and then on the solution data. In order to maintain the general ability of the model in the first stage, only the parameters of the query and value projection layers of the Transformer are updated instead of full-parameter training. c) Data augmentation: In this approach, we utilize the metaset as the augmented data of conventional instruction tuning methods that the metasets are mixed with solution data and the model is trained uniformly. The model trained on the mixed data is referred to as MetaTool.
@@ 3 Experiments
In this section, we evaluate our approach in both tool-oriented agent and tool-augmented chatbot scenarios. On the one hand, we fine-tune the LLM to master a specific toolset for achieving various complex objectives. On the other hand, we conduct large-scale instruction tuning to enable the model to generalize to new tasks and understand new tools through zero-shot documentation.
3.1 Tool-oriented Scenarios
3.1.1 Task setup
We adopt 3 tool-oriented tasks that emphasize complex tool execution and sequential planning. Among them SAW is newly designed while the other two are introduced from PlanBench Valmeekam et al. (2024). The key challenge of these tasks is to understand the rules (preconditions) and the environmental dynamics caused by actions. The task definition and dataset construction are elaborated below.
SpellAnyWord (SAW). In this task, the agent needs to sequentially construct a string that contains the target string as a continuous substring. The initial state of the task is a void string. Two non-degradable tools (functions) are avaliable: Add: to add two adjacent letters in the alphabet to the end of the current string. The tool input $\theta$ should be the preceding letter (e.g. passing ’a’ to Add on current string ” will result in ’ab’). Swap: to swap the position of two adjacent letters in the current string. The input should be the preceding letter (e.g. passing ’a’ to Swap on ’ab’ will result in ’ba’). An example task: The target string is ’any’. A successful action sequence can be [Add(’a’), Add(’n’), Add(’y’), Swap(’a’), Add(’o’)], which will result in a state sequence [’ab’, ’abno’, ’abnoyz’, ’banoyz’, ’banyoz’] and the final string ’banyoz’ has ’any’ as a substring. To eliminate the basis from tokenization, we format each string as a list of letters in prompts throughout the task.
@@ 3.1.1 Task setup
Datasets collection. For the SAW task, we randomly sample 2k target strings (from 2 letters to 10 letters) as task goals. We modify the BlocksWorld and Logistics tasks from PlanBench into the tool-use version, thus 2k goals for each task are adopted following the original configuration. Optimal action sequences are obtained with heuristic strategy as the solution data. For each annotated action, we generate a thought with GPT-generated templates. The thoughts analyze the situation and what to do following ReACT Yao et al. (2022) to leverage the model’s reasoning ability. Thus the solution contains a sequence of thought-tool-input tuples.
3.1.2 Implementation details
Our model is fine-tuned based on LLaMA3-8b-instruct AI@Meta (2024) with parameter-efficient fine-tuning method Qlora Dettmers et al. (2024) on 8 A100 GPUs. We utilize the instruction tuning version of LLaMA3 since comprehending tool-oriented tasks with specific objectives is the basis of tool understanding and use. For each task, we train the model on 10k meta-task data and 10k solution data for 3 epochs with AdamW optimizer and the learning rate of 2e-4. The models are tested in a simulated environment that receives the action of using a tool and returns the outcome and current state. We evaluated the model performance on 100 unseen cases of each three tasks.